### v2

# Correlation-induced phase shifts and time delays in resonance enhanced high harmonic generation from Cr+

Abstract
We investigate resonance-enhanced high harmonic generation (rHHG) in Cr+ by comparing a 1D shape-resonant model, time-dependent density functional theory (TDDFT), and independent particle approximation (IPA) simulations. Previous studies linked rHHG to the $3p\rightarrow 3d$ giant resonance, suggesting a modified four-step model where recolliding electrons are first captured in the autoionization state before recombining into the ground state, potentially leading to an emission delay associated with the resonance lifetime. While both the 1D-model and TDDFT reproduce experimental spectra, TDDFT reveals that rHHG counterintuitively originates from spin-down $3p$ states, while spin-up $3d$ electrons negligibly contribute. The IPA fails to reproduce rHHG, highlighting the significance of electron correlations. Furthermore, TDDFT revealed a correlation-induced $\sim$480 attosecond time delay, accompanied by a strong phase shift across the resonance, potentially explaining earlier RABBIT measurements. Our work sheds light on long-standing open questions in rHHG and should advance novel ultrafast spectroscopies of electron correlations and resonances.
High-harmonic generation (HHG) is a fundamental process in strong-field physics, enabling generation of coherent extreme ultraviolet (XUV) and soft X-ray radiation[1, 2, 3, 4, 5]. It presents a powerful tool for probing ultrafast electron dynamics and structural properties of matter across phases [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. In gases, HHG is well described by the three-step model [22, 23] whereby an electron: (i) tunnel ionizes, (ii) is accelerated by the laser, and (iii) recombines with the parent ion, emitting a photon. The coherence of the electronic wave function throughout this process is key towards utilizing HHG for ultrafast spectroscopy, as has been demonstrated for charge migration[16, 24], molecular symmetry and orientation [25, 26, 27, 28, 29, 30, 31, 32, 33], and more.
One known gas-phase system that deviates from this three-step model picture is transition-metal ions. Experimentally, HHG from laser-ablated plasma is known to lead to a significant resonant enhancement of a single harmonic order [34, 35, 36]. Resonant HHG (rHHG) has been established to originate at the single atom level, with macroscopic effects playing a negligible role [36]. The energy of the enhanced harmonic photon aligns with a transition to an autoionization state (AIS) in the cation, related to the $3p\rightarrow 3d$ giant resonance. Briefly, this autoionization transition involves the excitation of a $3p$ core electron to an unoccupied $3d$ subshell [37, 38, 39], after which the resulting highly excited state (e.g., $3p^{5}3d^{n+1}$) is unstable and decays via autoionization, typically filling the $3p$ hole and emitting an outer electron (e.g., $3p^{6}3d^{n-1}+\epsilon l$). Resonances in photoionization spectra and resonant photoionization have been the subject of extensive study. A complete topical review is provided in [40]. For Cr+, the ground state configuration is $3p^{6}3d^{5}(^{6}S)$ in LS coupling notation, corresponding to a half-filled $3d$ sub-shell with five spin-up electrons. This configuration breaks symmetry between spin-up/down electrons due to an exchange interaction. The dominant excitation is still the $3p\rightarrow 3d$ transition, i.e. Cr+ $3p^{6}3d^{5}(^{6}S)+\gamma$ → Cr+ $3p^{5}(^{2}P)3d^{6}(^{5}D)(^{6}P)$ followed by autoionization into Cr $^{2+}\penalty 10000\ 3p^{6}3d^{4}(^{5}D)+\epsilon l$ with $l=3$ being the dominant channel [37, 39, 38].
To account for enhancement in rHHG in presence of multi-photon transitions, an extended semi-classical single-active electron (SAE) 4-step model was suggested [41]. It replaces the recombination step of the three-step model with electron capture in the AIS, followed by a transition to the ground state and photon emission.
Alternatively, one can describe rHHG fully numerically with methods falling into two categories: (i) SAE models where the AIS is approximated by a resonance in a barrier-well potential, forming a metastable state (MS) [41, 42, 43, 44]. These models qualitatively reproduces the measured enhancement, but do so by synthetically fitting the transition energies. It remains unclear to what extent the SAE approximation captures the full dynamics of rHHG and if any time-delay or phase shift are inflicted on the resonant photon. Approach (ii) is to perform ab-initio simulations of rHHG using either time-dependent density functional theory (TDDFT) [45, 46], or multi-configuration time-dependent approaches [47, 48]. These schemes lead to spectra in agreement with experiments without the need of external parameters, but are more difficult to analyze and extract the physical mechanism. Still, with ab-initio schemes it was established that in some systems correlations play a dominant role in enabling resonant emission [45], and that deeper orbitals often play a crucial role in dynamics [47]. What remains unclear is how these effects transpire, and how they depend on the electronic spin in magnetic ions such as Cr+. In particular, the phase and time delay of rHHG has not yet been explored with multi-electron theories (only with model semi-analytic approaches[42]).
We perform here an extensive theoretical analysis of rHHG from magnetic Cr+ employing three complementary levels of theory: (i) a shape-resonant SAE model, (ii) ab-initio spin-TDDFT simulations, and (iii) simulations on par with the spin-TDDFT but where electron correlations are artificially frozen, denoted as the independent particle approximation (IPA). This comparison allows pin-pointing the mechanism of rHHG - a low-lying $3d$ electron is tunnel ionized and gains high kinetic energy in the laser field. When it re-collides with the parent ion, it promotes a deep-level $3p$ electron via electron-electron interactions into empty $d$-bands, forming the autoionization state, from which an enhanced rHHG photon is emitted. The $3p$ energy level is situated too deep to typically contribute to HHG if electronic interactions are not invoked. We show that this transition, which does not occur in uncorrelated theories, inflicts a time-delay and phase shift onto the rHHG emission, suggesting correlated corrections to the 4-step model. Importantly, we find that this transition is only facilitated by spin-down electrons that can pass through the empty $d$-bands. Our results shed light onto fundamental rHHG physics and could lead to HHG-spectroscopy of attosecond correlated multi-electron systems.
Let us briefly describe our numerical approach and examined set-up. First, we explore HHG and corresponding time-frequency plots from a 1D SAE model that includes a shape resonance matching withe the experimental resonant transition (all numerical details are delegated to the Appendix). The results in Fig. 1 show a good agreement with experimental rHHG spectra from Cr+ [36, 49]. We examine the behavior of resonant H29 for different resonance lifetimes (Fig. 1 b), especially quantifying its enhancement factor, defined as $\frac{I_{29}}{I_{27}}$. In doing so, we hope to resolve whether a re-collision picture is valid in rHHG in the SAE case. While the overall harmonic yield decreases with increasing lifetime (see SI), the enhancement factor increases. This is consistent with a recollision-based picture, since the increase in lifetime is achieved by widening the barrier, thereby decreasing the tunneling rate. To further investigate this interpretation, we gradually extended the complex absorbing potential (CAP) inwards from the box boundary toward its center. As the CAP approached the semiclassical long trajectories, these progressively disappeared from the time-frequency (Fig. S2), while short trajectories were unaffected. By extending the CAP inward the harmonic yields reduce, including H29, further supporting the recollision mechanism for rHHG.
Time-frequency analysis for an exemplary case is shown in Fig. 1(c). The semiclassical trajectories match the quantum SAE simulations with high accuracy — there is no distinguishable time delay for the resonant H29. We note that according to the 4-step model, H29 is emitted after the electron is captured in the AIS, therefore some delay is anticipated, though we could not resolve it in our conditions.
Figure 1: rHHG from the 1D model. (a) Potential supporting a shape resonance (MS, black). The ground state (blue) and MS (green) densities are plotted in their respective eigen-energies. The vertical axis (energy) is normalized to the driving photon energy. (b) HHG spectra for different barrier widths/MS lifetimes. Curves are normalized to the 27th harmonic. The enhancement coefficient increases with the MS lifetime. (c) (Right) Gabor transform of HHG emission for an MS lifetime of 0.17T (T is the period of the driving laser). Red, purple, and black curves indicate the resonant harmonic energy, the semiclassical cutoff energy, and semiclassical emission times (trajectories), respectively. (Left) Corresponding spectrum. The enhancement of the resonant harmonic H29 is evident in both the spectrum and the Gabor transform. The harmonic emission times are in strong agreement with the semiclassical model.
Next, we employ an ab-initio TDDFT scheme (for details see Appendix). Figure 2 shows HHG spectra and time-frequency analysis under similar conditions to Fig. 1 separated to spin-up/down channels (Fig. 2(c,d)), and the total spin-summed response (Fig. 2(b)). While both channels contribute to the resonance emission, the spin-down channel response is about an order of magnitude larger than the spin-up, dominating rHHG. This is counter-intuitive if considering the typical physical mechanism of atomic HHG and the energy level diagram in Fig. 2(b) - the spin-up channel includes initially occupied $3d$ states that have a much lower ionization potential compared to the $3p$ levels. A-priori, this should make their tunnel ionization and recombination exponentially more efficient. However, the fact that $3p$ spin-down states dominate the resonant response indicates that they become intrinsically linked to the AIS through interaction with the laser. In these calculations, the long trajectories are slightly suppressed compared to the 1D case, as is common in ab-initio theory[50], due to the 3D nature of the simulation.
Another surprising feature can be seen in the time-frequency plots in Fig. 2. First, we notice that the HHG emission time in the spin-up channel largely follows the semi-classical trajectories with ionization potential matching the $3d$ levels. This is expected since the $3d$ spin-up sub-shell is initially occupied and its ionization energy is much lower than the ionization energy of the $3p$ electrons. The most striking feature is that the HHG emission from the spin-down channel also largely follows the classical trajectories initiated from $3d$ electrons, even though the $3d$ states are not initially occupied by spin-down electrons. This is a direct indication that rHHG induces a $3p\rightarrow 3d$ transition, while this channel is blocked for the $3p$ spin-up electrons, as the $3d$ spin-up bands are fully occupied.
[…]
Figure 2: rHHG with TDDFT. (a) Energy level configuration in Cr+. (b) (Right) Gabor transform of total calculated HHG emission. (Left) Corresponding HHG spectrum. Top and bottom red curves denote the emission time and energy according to the classical trajectories of the 3-step model (low energy curves associated to $3d$ electrons and high energy curves are associated to $3p$ electrons). White lines illustrate the driving field. (c) Same as (a) but for spin-up response, roughly following the semiclassical model for $3d$ electrons. (d) Same as (c) but for spin-down response, where $3p$ spin-down states dominate rHHG (see color scales in (c) and (d)), which is delayed by 250 as relative to spin-up response.
To further explore this conjecture, we invoke the IPA, i.e. freezing electronic interactions altogether. The results are shown in fig. 3 for both spin channels and the total emission, and reflect a prominent effect - upon freezing correlations the resonance enhancement is completely suppressed. Similar results were recently shown in another system[45]. Here however, the spin-up channel HHG response remains largely similar with/without the IPA, while the spin-down HHG response is substantially suppressed and is the source of the disappearance of rHHG. Notably, the emission from the spin-down electrons no longer follows the semi-classical trajectories of the $3d$ electrons, proving that the spin-down $3p\rightarrow 3d$ transition is facilitated by dynamical electron correlations even in the tunneling regime. Further, while the resonance emission line still exists in the IPA spin-down channel, it is orders of magnitude weaker, has phase shifted, and is now aligned with the maximum of the electric field, similar to the perturbative low-order harmonics. In that respect, even though the atomic system still supports a resonant state in the absence of correlations, it does not contribute to rHHG since emission from typical HHG channels is more prominent, and the resonance emission mechanism differs from that of the correlated system.
The spin-down emission spectra in both the IPA and correlated models closely resemble the 1D model, exhibiting a dip at lower harmonics followed by a peak at resonance. In the 1D case, this has been attributed to the returning electron’s reduced ability to penetrate the potential barrier [43].
Overall, we conclude that in the IPA, the $3p$ down states do not transition into $3d$ states, and the resonance is emitted by pure multi-photon transitions rather than excitation through the continuum. In contrast, spin-down $3p$ electrons require electronic correlations in order to resonantly emit HHG in Cr+ in experimental settings, following their transition into $3d$ levels, also incurring a time delay. These results are further corroborated by ab-initio simulations for the single-photon absorption and ionization cross sections (see SI), supporting our conclusions that the down spin channel is the main contributor to absorption at the resonance.
Figure 3: rHHG within the IPA. (a-c) Same as (b-d) in Fig.2. Note the spin-down response on resonance is in-phase with the maxima/minima of the driving field and strongly suppressed, unlike in Fig.2.
Finally, we explore the phase of the harmonics at all levels of theory with a RABBIT-like simulation. We extract the harmonic phase across orders, as well as the RABBIT sideband (SB) visibility per order (see SI for details). The results presented in Fig. 4(a) show a sharp negative phase jump across the resonance that arises only in the correlated theory. This shift is up to $2.3\pi$ through the resonance, accompanied by a reduction in SB visibility. A vertically shifted IPA curve which aligns with the start and end of the correlated curve shows excellent agreement between the models, except near the resonance. We have found that differences in the 1D and IPA simulations can be corrected for with group velocity delay corrections (blue line in Fig. 4(a), see SI for details), rendering both theories analogous. The phase jump in rHHG can potentially explain previous measurements of reduced SB visibility [51], since it suggests a sharp phase jump even within the resonant harmonic itself (the $2.3\pi$ shift occurs abruptly over 3 harmonic orders).
To further test this, Fig. 4(b) presents the actual numerically-observed RABBIT SB visibility and the expected visibility due to two-source interference with unequal intensities. The normalized visibility is shown in Fig. 4(c). For both SB28 and SB30 (adjacent to the resonance), across all models, the expected visibility is larger than the observed with the exception of SB30 IPA spin-down, where they are roughly equal. For both SBs, the normalized correlated visibility is drastically lower than the normalized IPA visibility (SB28 - 15.1% vs 27.1%; SB30 5.6% vs 35%). The spin-resolved components show similar behavior (most noticeably the spin-down component in SB30 (4% vs 103.5%)). An outlier is the spin-up component of SB30 (48.5% vs 4.5%), which we attribute to the extremely low harmonic signal in this channel. Interestingly, the 1D SAE model yields a normalized visibility similar to that of the spin-down component in the correlated model. This contrasts with the harmonic phase, which more closely resembles that of the IPA model. While this may seem unexpected, it is important to note that the 1D and IPA models are not equivalent. The 1D model employs a synthetic potential supporting a sharp shape resonance, which may differ from the IPA simulations, where this resonance could be less pronounced. Crucially, only the fully correlated theory accurately reproduces the correct harmonic phase near resonance, also strongly reducing SB visibility in correspondence with measurements[51].
Figure 4: rHHG phase and RABBIT analysis. (a) Harmonic phase comparison with varying levels of theory. Yellow and orange curves denote the fully correlated and IPA simulations, respectively. The blue line denotes the group-delay-corrected 1D simulation. Gray dotted lines with dot markers indicate the vertically shifted IPA, while the gray dotted line with $\times$ markers corresponds to the unadjusted 1D result. (b) Observed and expected SB visibility (for SB28 and SB30) for various levels of theory and spin channels. (c) Normalized visibility for both sidebands and various levels of theory. The correlated theory shows strongly reduced normalized visibility, corresponding to the strong phase jump in (a).
In summary, we theoretically explored resonance-enhanced high harmonic generation (rHHG) in Cr+ ions by employing three complementary approaches: (i) a 1D SAE model where the resonance is modeled by a shape-resonance, (ii) ab-initio spin-time-dependent density functional theory (spin-TDDFT), and (iii) spin-TDDFT within the independent particle approximation (IPA) with electronic interactions artificially removed.
The 1D model successfully reproduces experimental rHHG spectra, where we showed that the mechanism for rHHG in this case is tunnel-ionization followed by re-collision (without associated time delays). TDDFT also reproduces experimental spectra, but offers a more comprehensive view into rHHG physics — we uncovered that the primary contribution to enhancement originates from spin-down $3p$ states rather than spin-up $3d$ electrons (that intuitively should dominate due to their lower ionization potential). Strikingly, the more dominant spin-down emission is delayed by $\sim$480 attoseconds compared to the spin-up $3d$ emission. These effects contrast with the 1D model and IPA simulations and point to a mechanism involving electron interaction and spin selectivity, also aligning with the 4-step picture: a $3d$ spin-up electron tunnel ionizes and accelerates in the continuum, re-collides with the parent ion, exciting a $3p$ spin-down electron (only possible through correlations), exciting an autoionization state. This proceeds to coherently decay to the ground state, emitting a resonantly-enhanced photon. Lastly, we explored RABBIT simulations that investigate the harmonic phases and sideband visibility. These simulations showed a sharp phase jump ($\sim 2.3\pi$) across the resonance, only in the correlated theory. Correspondingly, normalized sideband visibility near the resonance was drastically lower in the correlated theory compared to the IPA (especially for the dominant spin-down component), which might explain earlier measurements[51].
Overall, our work establishes that the resonance enhancement mechanism is physically different in the absence of correlations, and thus suggests that SAE and semi-classical models must be phase and time-delay corrected due to electronic interactions. Our work resolves several open questions in the field, and should pave way to novel ultrafast spectroscopies of electron correlations and resonances.
Acknowledgments. ON gratefully acknowledges the scientific support of Prof. Dr. Angel Rubio and the Young Faculty Award from the National Quantum Science and Technology program of Israel’s Council of Higher Education Planning and Budgeting Committee. GM acknowledges support from the Israel Science Foundation Grant No. 65618 and Grant No. 89223. AP acknowledges from the Israel Science Foundation Grant No. 1484/24 and the Alon Fellowship of the Israeli Council of Higher Education.